OpenStudy (anonymous):

if f(x)=1/3x-1 and g(x) = f^-1(x), find g’(x).

OpenStudy (bibby):

we literally just did this?

OpenStudy (lexi724):

g o f) (x) = 2 ( 3x^2 - 1) = 6x^2 - 2 two: An = 5 * 2^n three can't do with text

OpenStudy (anonymous):

Thats incorrect darling

OpenStudy (bibby):

Don't patronize her

OpenStudy (anonymous):

Please reference this question http://openstudy.com/study#/updates/52c33e98e4b0b729fb8c0a95

OpenStudy (anonymous):

is it \[f(x)=\frac{1}{3x-1}\]??

OpenStudy (bibby):

whoops

OpenStudy (anonymous):

@satellite73 yes

OpenStudy (anonymous):

one method is to actually find the inverse, and then take the derivative of that

OpenStudy (anonymous):

not too hard to find \[x=\frac{1}{3y-1}\] \[3y-1=\frac{1}{x}\] \[3y=\frac{1}{x}+1\] \[y=\frac{1}{3x}+\frac{1}{3}\] or \[f^{-1}(x)=\frac{1}{3x}+\frac{1}{3}\] now the derivative is easy to find

OpenStudy (anonymous):

i mean i guess it is easy to find, right?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

oh heck no

OpenStudy (anonymous):

grrrr

OpenStudy (anonymous):

the derivative of \(\frac{1}{x}\) is \(-\frac{1}{x^2}\)

OpenStudy (anonymous):

so the derivative of \(\frac{1}{3}\) is \(-\frac{1}{3x^2}\)

OpenStudy (anonymous):

oops typo there, i mean the derivative of \[\frac{1}{3x}\] is \[-\frac{1}{3x^2}\]